PROPERTIES OF NEARLY ONE-ELECTRON MOLECULES: A GREEN’S FUNCTION APPROACH

\underline{SERHAN N. ALTUNATA}, STEPHEN L. COY, ROBERT W. FIELD, Department of Chemistry and G.B. Harrison Spectroscopy Laboratory, Massachusetts Institute of Technology, Cambridge, MA 02139.

We present an improved procedure to determine the collision theory reaction matrix (K matrix) and bound state energy levels in an effective one-electron calculation. As a first illustration, we consider the scattering of an electron from a monopole and a point dipole [J. K. G. Watson, Mol. Phys. 81, 277 (1994).] and obtain excellent agreement with the exact results. In the second illustration, the spectral quantum defects for the ^{2}Σ^{+} states of CaF are computed using the effective potential of [M. Arif, Ch. Jungen, and A. L. Roche, J. Chem. Phys. 106 4102 (1997)], obtaining results close to the experimental data. We show that other representations of the ion-core are compatible with our treatment.